OpenStudy (anonymous):

f(x) = 3 x^{6} \sqrt{x} + \frac{-3} {x^{2} \sqrt{x} }. how do you find the derivative

geerky42 (geerky42):

\[f(x) = 3 x^{6} \sqrt{x} + \frac{-3} {x^{2} \sqrt{x} }\]?

geerky42 (geerky42):

Just apply sum and power rule.

OpenStudy (anonymous):

yes but i dont know how to do this

OpenStudy (anonymous):

okay so use the product rule for the first half and then use the quotient rule for the second half

geerky42 (geerky42):

\(x^n\sqrt x = x^{n+1/2}\)

geerky42 (geerky42):

Nah, we only need power rule

geerky42 (geerky42):

So... \[\begin{align} 3 x^{6} \sqrt{x} + \frac{-3} {x^{2} \sqrt{x} } &= 3 x^{13/2} - \frac{3} {x^{5/2}}\\~\\&= 3 x^{13/2} - 3x^{-5/2} \end{align}\] Now apply power rule: \(\dfrac{d}{dx}x^n = nx^{n-1}\)

OpenStudy (anonymous):

what is n

geerky42 (geerky42):

Any number

geerky42 (geerky42):

While taking derivative of x, you just multiply it by its exponent, then subtract its exponent by 1.

OpenStudy (anonymous):

yes but if i put that as my answer it comes out wrong

geerky42 (geerky42):

You don't put \(d/dx~x^n = nx^{n-1}\) as answer... It's just power rule you need to apply.

OpenStudy (anonymous):

so how can i find the answer

geerky42 (geerky42):

Let's do a few easy questions. What is derivative of \(x^2\)?

OpenStudy (anonymous):

2x

geerky42 (geerky42):

Yes, what about \(x^3\)?

OpenStudy (anonymous):

i know how to apply the power rule

geerky42 (geerky42):

ok, what are you struggle with?

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

to find the answer

geerky42 (geerky42):

Well, just take derivative, then simplify

geerky42 (geerky42):

Remember sum rule; basically you just derivative each term individually.

OpenStudy (anonymous):

so its 13/3 times 3

geerky42 (geerky42):

13/2 times 3. Times x with what exponent?

OpenStudy (anonymous):

i dont follow

geerky42 (geerky42):

\(3*\dfrac{13}{2}x^?\)

OpenStudy (anonymous):

dont you subtract 13/2-1

geerky42 (geerky42):

Yeah, so \(3*\dfrac{13}{2}x^{11/2}\) Now do next term. can you find derivative of that term?

OpenStudy (anonymous):

7.5/x^7/2?

geerky42 (geerky42):

Yeah, so you have \(\dfrac{d}{dx}f(x) = \boxed{\dfrac{39}{2}x^{11/2} + \dfrac{15}{2x^{7/2}}}\)

OpenStudy (anonymous):

how did you get 15

geerky42 (geerky42):

\(7.5 = \dfrac{15}{2}\)

OpenStudy (anonymous):

Thank you so much I appreciate it

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

how did you get x^5/2 in the begininng

OpenStudy (anonymous):

and 3x^13/2

geerky42 (geerky42):

\(x^2\sqrt x = x^{2+\frac{1}{2}}=x^{\frac{4}{2}+\frac{1}{2}} = x^{5/2}\) Same for \(x^{13/2}\)

OpenStudy (anonymous):

Thank you I get it